### v1

# Temperature sensitivity of a Thorium-229 solid-state nuclear clock

Abstract
Quantum state-resolved spectroscopy of the low energy thorium-229 nuclear transition was recently achieved. The five allowed transitions within the electric quadrupole splitting structure were measured to the kilohertz level in a calcium fluoride host crystal, opening the field of nuclear-based optical clocks. Central to the performance of solid-state clock operation is an understanding of systematic shifts such as the temperature dependence of the clock transitions. In this work, we measure the four strongest transitions of thorium-229 in the same crystal at three temperature values: 150 K, 229 K, and 293 K. We find shifts of the unsplit frequency and the electric quadrupole splittings, corresponding to decreases in the electron density, electric field gradient, and field gradient asymmetry at the nucleus as temperature increases. The m = $\pm$5/2 → $\pm$3/2 line shifts only 62(6) kHz over the temperature range, i.e., approximately 0.4 kHz/K, representing a promising candidate for a future solid-state optical clock. Achieving 10-18 precision requires crystal temperature stability of 5 $\upmu$K.
preprint: APS/123-QED
Introduction. Nuclear clocks promise a new regime of scientific measurement capability. This includes searches for ultralight dark matter candidates through variations of fundamental constants [1, 2, 3, 4], solid-state optical clocks for field applications [5, 6, 7], and more precise clock operation in trapped ions [8, 9, 10]. Current optical atomic clocks based on electronic transitions utilize quantum state engineering for precise measurement of high quality factor resonances and detailed characterization of systematic effects to account for or reduce shifts and broadenings [11, 12, 13]. Examples in these cold atom systems include blackbody radiation shifts, light shifts, and Zeeman splittings and shifts [14]. To achieve the level of precision of optical atomic clocks, nuclear clock transition shifts will need to be similarly characterized.
Figure 1: (a) Schematic of the crystal mount and thermal control. The thorium-doped calcium fluoride crystal is mounted on a baseplate in vacuum, which is connected to an external dewar via a copper thermal link. The temperature is continually monitored using a sensor attached to the mounting plate. The dewar is in contact with a cold trap which condenses molecules to reduce vacuum pressure. The dewar is filled with liquid nitrogen for measurements at 150 K, a methanol/dry ice mixture for measurements at 229 K, and is unfilled for measurements at 293 K. A parabolic mirror collimates fluorescent photons, which are then counted downstream after spectral filtering. Shown in the inset is the crystal structure of calcium fluoride, where one calcium ion is replaced with doped Th4+, and two F interstitials are added for charge compensation. (b) Frequency comb spectroscopy of thorium. The comb mode number exciting the transition was determined in previous work (see text) [3]. The frequency of each spectral line is determined by scanning frep and fitting to the line center. (c) Electric quadrupole structure of 229Th. Increasing the electric field gradient causes the lines to spread outward as splitting increases. Increasing the electron density at the nucleus induces an absolute shift of all lines. Not shown is the m = $\pm$1/2 → $\pm$3/2 transition, whose intensity is an order of magnitude weaker than the other four lines and is not included in this study.
The thorium-229 nucleus has a nuclear transition (229mTh) at 8.4 eV (2020 THz), accessible with vacuum ultraviolet (VUV) lasers [15, 16, 17, 18]. The last two years have seen rapid progress in the development of nuclear-based clocks [19, 20]. Most recently, a coherent VUV frequency comb was used to resolve individual quantum states of the nuclear transition in a calcium fluoride (CaF2) crystal, and connected the absolute frequency to the 87Sr optical atomic clock as an absolute frequency reference [3]. This new nuclear frequency standard placed the measurement uncertainty of the 229mTh transition at the kilohertz level. The resolved nuclear electric quadrupole structure allowed the determination of the ratio of the ground state’s electric quadrupole moment to that of the isomeric state to be $Q_{\rm e}$/$Q_{\rm g}$=0.57003(1).
The thorium nuclear transition is an excellent candidate for a next-generation optical frequency standard [21, 5, 8, 22]. The 229mTh lifetime in CaF2 was measured to be 641(4) s, placing the projected oscillator quality factor at potentially the 1019 level. In addition, the small multipole moments of nuclei compared to electron orbits make the nuclear transition much less sensitive to external perturbations. It has been suggested that the transition will have an enhanced sensitivity to variations of fundamental constants such as the fine structure constant $\alpha$ [1, 5]. Using the quadrupole ratio measured in Ref. [3], the enhancement factor was estimated to be three orders of magnitude higher than that in current optical atomic clocks [4, 23].
Besides the application in metrology, the thorium nuclear transition can also be used to precisely probe material properties such as the local electron density. Mössbauer spectroscopy uses nuclear transitions to study the chemical environment surrounding the nucleus, including oxidation state, local magnetic field couplings, and electron distribution [24, 25]. It is useful for many applications such as spin crossover in organometallic compounds [26], bioinorganic systems [27] including metalloproteins [28], nanomaterials [29], and temperature dependent studies of solids [30]. There are over 45 active Mössbauer elements used today [25], including thorium-232 [31]. Mössbauer spectroscopy is typically conducted using gamma ray sources or synchrotron radiation. The thorium-229 nucleus is the first and possibly only Mössbauer active transition accessible by current tabletop laser sources.
In this work, we measure the 229mTh transition at multiple temperature (T) values and extract the T-dependent electric quadrupole splittings and the chemical isomer shift. We confirm that the nuclear quadrupole ratio remains constant while the extracted electric field gradient (EFG) and the absolute transition frequency show clear changes with respect to temperature. These indicate that as T increases, the electron density and EFG at the nucleus both decrease (see discussion below). One line, the m = $\pm$5/2 → $\pm$3/2 transition, is particularly insensitive to temperature change due to cancellation of the splitting from the absolute shift, making it suitable for future stable clock operation.
Figure 2: Frequencies of the four prominent electric quadrupole lines plotted against temperature. Shown at the top are zoomed views of the full spectrum, plotted at the bottom. The frequency axes of the top plots all have a 3 MHz range. The red line shows the absolute shift, representing the unsplit I = 5/2 → 3/2 transition. The absolute frequency decreases with increasing temperature, indicating that electron density at the nucleus decreases. The four lines show reduced quadrupole splitting when temperature is increased, indicating a smaller electric field gradient. The m = $\pm$5/2 → $\pm$3/2 transition is the least temperature sensitive transition due to a cancellation of these two effects.
Methods. We measure the absolute frequency of the transition through direct VUV frequency comb spectroscopy, described in detail in previous publications [3, 32, 33]. Our setup is shown in Fig. 1a. A 229Th-doped calcium fluoride crystal (229Th:CaF2 grown at TU Wien, see Refs. [34, 35] for details) is cooled in a vacuum chamber via an external dewar connected to a cold trap and a copper thermal link. The reported T is read out using a sensor attached to the same mount as the crystal. Measurements were done at three temperatures: 150(1) K using liquid nitrogen, 229(1) K using a mixture of dry ice and methanol, and 293(1) K at room temperature. After initial equilibration, T remains stable to within 1 K for each T range during the measurement time. We conduct scans by changing the come repetition frequency, which shifts the VUV comb mode spacing.
For each spectroscopic feature, the comb mode exciting the transition at 150 K was determined in our previous work [3]. We expect that the T-dependent shift is on the order of $\sim$1 MHz or less [5, 6], which is much less than the comb mode spacing of 75 MHz (Fig. 1b). Thus, the comb mode number exciting a given transition remains the same for all temperatures. We find that the measured transition lines are in strong agreement with expected electric quadrupole splittings (see below), further indicating that our assumption is valid. The center frequency of each line is determined using previously described methods [3]. The absolute frequency for each feature is determined from the frequency comb equation,
$\nu_{\mathrm{Th}}=Nf_{\mathrm{rep}}+7f_{\mathrm{CEO}},$ (1)
where $N$ is the integer comb mode number ($\sim$2.7 $\times$ 107), $f_{\rm rep}$ is the comb repetition frequency, and $f_{\mathrm{CEO}}$, the carrier envelope offset frequency, is –8 MHz in this study. The uncertainty of each line is determined by multiplying the center frequency uncertainty in frep by the comb mode number.
Absolute line frequencies. The 229Th ground state has a nuclear spin of I = 5/2, while the isomeric state 229mTh has nuclear spin I = 3/2. This gives rise to five allowed magnetic dipole transitions given by the selection rule $\Delta m$ = 0 or $\pm$1, where $m$ is the angular momentum projection onto the z-axis. We scanned four of the five allowed transitions within the electric quadrupole structure. The absolute frequency for each spectroscopic line at the three temperatures is shown in Table 1.
The fifth line corresponding to m = $\pm$1/2 → $\pm$3/2 was not scanned due to its relatively weaker transition strength [6] but is calculated using the sum rule described in Ref. [3]. These calculated transitions are listed in Table 1 in gray. Fig. 2 shows the measured transition energies plotted against T, shown in black. The top section is a series of zoomed in plots with the frequency axis spanning 3 MHz for each line to show the relative T-dependence of the four features.
Plotted in red is the absolute, i.e. unsplit, transition frequency, calculated using the formula
$\nu_{\rm Th,abs}=\frac{1}{6}(\nu_{\frac{3}{2}\rightarrow\frac{1}{2}}+2\nu_{\frac{5}{2}\rightarrow\frac{3}{2}}+2\nu_{\frac{1}{2}\rightarrow\frac{1}{2}}+\nu_{\frac{3}{2}\rightarrow\frac{3}{2}}).$ (2)
The absolute transition frequency, free from quadrupole splittings, is calculated by taking a proper average of the four measured transitions; this calculation is valid even for a non-zero asymmetry parameter $\eta$. The absolute frequency is reduced by 776 kHz at room temperature relative to 150 K, indicating a change in electronic charge density at the nucleus. In addition, the peak splittings decrease as T is increased, due to the change in EFG. Three of the peaks exhibit a clear shift on the order of a few MHz. One peak, the m = $\pm$5/2 → $\pm$3/2 transition, shows roughly one order of magnitude smaller sensitivity to temperature, with a shift of only 62(6) kHz over the 143 K range of temperature changes.
Table 1: Measured (black) and calculated (gray, starred) frequencies at three different temperatures, corresponding to Fig 2. The EFG-free field is calculated using Eq. 2.
Transition
Abs freq (MHz)
150(1) K
Abs freq (MHz)
229(1) K
Abs freq (MHz)
293(1) K
(A) 3/2 $\rightarrow$ 1/2 2020 407 283.847(4) 2020 407 284.808(8) 2020 407 285.662(5)
(B) 5/2 $\rightarrow$ 3/2 2020 407 298.727(4) 2020 407 298.722(3) 2020 407 298.784(5)
(C) 1/2 $\rightarrow$ 1/2 2020 407 446.895(4) 2020 407 445.654(5) 2020 407 444.551(3)
(D) 3/2 $\rightarrow$ 3/2 2020 407 530.918(4) 2020 407 529.996(5) 2020 407 529.021(3)
(E) 1/2 $\rightarrow$ 3/2 2020 407 693.98(2) 2020 407 690.84(1)* 2020 407 687.910(7)*
EFG-free 2020 407 384.335(2)* 2020 407 383.926(3)* 2020 407 383.559(2)*
Temperature-dependent isomeric shift. There is an absolute, or chemical, shift that arises from nonzero electronic wavefunction density at the position of the nucleus [36, 37]. The electron density changes with T due to expansion of the lattice and a rearrangement of the electronic structure in response. The different charge radii of ground and excited nuclear states changes the electrostatic interaction between the electron density and the nucleus, inducing a shift in the nuclear transition energy (Fig. 1c). The thorium-229 nuclear transition is therefore dependent on the chemical environment of the host crystal. It is typically assumed that the primary contribution to the isomeric shift arises from s electrons due to their high wavefunction probability at the nucleus, although p, d, and f electrons can contribute to the shift through screening effects [37]. Further, T-dependent isomeric shifts are likely due to valence s electrons, as they participate in chemical bonding and are thus subject to environmental changes.
The T-dependent isomeric shift is given by [37],
$\delta\nu=\left(\frac{4\pi Ze^{2}R^{2}}{5h}\right)S^{\prime}(Z)[\delta\psi^{2}(0)]\left[\frac{\Delta R}{R}\right].$ (3)
Here, $\nu$ is the EFG-free transition energy, Z is the charge number of the nucleus, e is the electron charge, R is the nuclear radius, and $h$ is Planck’s constant. $S^{\prime}(Z)$ is a dimensionless “relativity factor” arising from relativistic corrections to electron charge density in heavy nuclei. In this work we use the value listed in Ref. [37] of $S^{\prime}(Z)$ = 11.68 for Z = 90. The factor [$\delta$$\psi$2(0)] is the change in the electron density at the nucleus (with units of inverse volume) as the crystal expands with T increase. The value [$\Delta$R/R] is the nuclear factor containing the change in nuclear charge radius $\Delta$R between the ground and isomeric states.
The reduction of the absolute transition frequency from 150 K to 293 K indicates that the electronic charge density at the nucleus decreases when T is increased. Using the value of $\Delta R^{2}$ = 0.0105(3) fm2 from Ref. [4] and the formula $\Delta R^{2}=2R\Delta R$, the change in electron density at the nucleus is [$\delta$$\psi$2(0)] = –0.161(5) Å-3. The quoted uncertainty does not account for approximations in calculating the relativity factor $S^{\prime}(Z)$.
[…]
Table 2: Fitted parameters at three different temperatures, corresponding to Fig 3.
[…]
150(1) 339.258(7) 193.387(5) 0.59163(5) 0
229(1) 337.320(4) 192.282(3) 0.58067(3) -409(4)
293(1) 335.331(9) 191.140(5) 0.57184(5) -776(3)
Temperature-dependent quadrupole splittings. The nuclear quadrupole structure arises from the interaction of the EFG of the crystal and the electric quadrupole moment of the nucleus. The Hamiltonian is given by [38]
$H=\frac{QV_{zz}}{4I(2I-1)}\left[3I_{z}^{2}-\textbf{I}^{2}+\eta\left(I_{x}^{2}-I_{y}^{2}\right)\right].$ (4)
Here, Q is the spectroscopic quadrupole moment in the laboratory frame with $Q_{\rm g}$($Q_{\rm e}$) being the ground state (isomeric) quadrupole moment, I is the total nuclear spin, and $\eta$ is the asymmetry parameter given by $\eta$ = ($V_{\rm xx}$-$V_{\rm yy}$)/$V_{\rm zz}$. In this formalism, we assume the axis where the EFG is diagonalized, and $V_{\rm zz}$$>$ $V_{\rm xx}$$>$ $V_{\rm yy}$ denote the x, y, and z components of the EFG. Nonzero values for $\eta$ induce state mixing between the quadrupole levels and therefore alter transition strengths. Our analysis accounts for the frequency shifts but does not use line intensity.
[…]
Discussion. In solid state nuclear clocks, the temperature dependence is a crucial systematic to characterize potential clock performance. Here, we measure four of the five transition lines within the electric quadrupole structure in thorium-doped calcium fluoride at three temperatures ranging over $\sim$150 K. We find that the frequencies of three of the transitions shift over a range of a few MHz when heated from 150 K to 293 K, representing a fractional shift of 10–12 K–1 to 10–11 K–1, in line with the T-dependent shifts predicted previously [5, 6]. Our measurement uncertainty at the kHz level is likely limited by the $\sim$300 kHz linewidth of the VUV frequency comb. We do not yet know the degree of inhomogeneous broadening due to the crystal environment, but this can be measured by further narrowing the comb linewith. Regardless, this study shows that the transitions have a maximum inhomogeneous linewidth on the order of $\sim$100 kHz, which is orders of magnitude lower than the majority of optical transition linewidths in the condensed phase [41, 42, 43, 44].
This study achieves the first observation of isomeric shifts using coherent lasers interrogating nuclear transitions. We see a reduction of the absolute transition frequency and a decrease in electric quadrupole splittings when T is increased, corresponding to a decrease in electronic charge density, EFG, and the asymmetry parameter at the position of the nucleus. Upon heating, electron density is pulled away from the thorium nucleus from expansion of the positions of nuclei, likely via the highly electronegative fluoride anions in the crystal lattice and/or in charge compensation sites. A deeper understanding of the microscopic changes in the crystal should be carried out using density functional theory or other electronic structure methods using these measured values as a benchmark. Electronic structure calculations that precisely calculate the change in charge density could possibly be used in tandem with Eq. 3 to extract more precise nuclear parameters such as the change in the nuclear charge radius [$\Delta R/R$] between the ground and isomeric state. This study opens the door for coherent Mössbauer spectroscopy, allowing for studies of these electronic structure shifts in other thorium-containing systems [24].
We observe one transition, m = $\pm$5/2 → $\pm$3/2, that is much less sensitive to temperature shifts – likely due to a fortuitous cancellation of the electric quadrupole splitting and absolute frequency shift. This transition has the strongest Clebsch-Gordan coefficient among the electric quadrupole splittings [6] and only changes by 62(6) kHz over the 143 K range, whereas the other transitions each change by more than 1 MHz. It is therefore a promising candidate for future clock operation. Precision at the 10-18 level corresponds to a frequency stability of $\sim$2 mHz, which requires stabilizing the crystal to 5 $\upmu$K, a level of temperature sensitivity demonstrated in multiple platforms [45, 46, 47]. We note that between 150 K and 229 K, the transition only changes by 5(5) kHz, indicating a possible further reduction in temperature sensitivity over this range; more data are required to verify this potential trend. Further, material engineering of the sample such as introducing controlled mechanical strain may offer a tunable method to change the crystal properties, possibly leading to further cancellation of temperature shifts and a more stable clock transition. Other systematic shifts and broadening mechanisms include magnetic dipole broadening (estimated to be on the order of $\sim$400 Hz) and second order Doppler effects (estimated shift at $\sim$1 Hz/K with broadening on the order of $\sim$1 Hz/K) [5, 6]. Future studies should measure and fully characterize these other systematics after narrowing the VUV comb linewidth [48].
Acknowledgements. We thank Kim Hagen for technical assistance. We also acknowledge funding support from the Army Research Office (W911NF2010182); Air Force Office of Scientific Research (FA9550-19-1-0148); National Science Foundation QLCI OMA-2016244; DOE quantum center of Quantum System Accelerator; National Science Foundation PHY-2317149; and National Institute of Standards and Technology. Part of this work has been funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 856415) and the Austrian Science Fund (FWF) [Grant DOI: 10.55776/F1004, 10.55776/J4834, 10.55776/ PIN9526523]. The project 23FUN03 HIOC [Grant DOI: 10.13039/100019599] has received funding from the European Partnership on Metrology, co-financed from the European Union’s Horizon Europe Research and Innovation Program and by the Participating States. K.B. acknowledges support from the Schweizerischer Nationalfonds (SNF), fund 514788 “Wavefunction engineering for controlled nuclear decays.”